[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Oct 04, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Oct 4 13:59:58 UTC 2017


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
1) TAFB did not have any ecmwf0p25 data from the 12z run (ecmwf0p25,  
ecmwf0p25_MSLP, ecmwf0p25_wave) in AWIPS. Only a few random timesteps for  
Wind and MSLP were available, with no timesteps for waves. Working theory  
is that it may be related to the NCEP network problems yesterday, so TSB is  
continuing to monitor this as we ingest the ECMWF manually into AWIPS  
throughout the day. Chris Mello turned on SBN 1.0 deg data until NCO fixes  
their network problems. This will show up in GFE as "ECMWF"

2) TAFB noticed that nhcn-ls3 is now prompting for a password for the ldad  
account, which it did not do before. TSB will investigate.

3) Matt Sardi informed TAFB that he and Jeff updated the TAFB sea state  
graphics to mimic the wind/waves by plotting gridded significant wave  
height values in addition to swell direction. These updates should improve  
efficiency due to less time spent adding labels across the entire chart and  
a consistent product with the format of the wind/waves.

4) TSB is continuing to look into the GOES imagery retrieval code, and will  
provide TAFB with proper instructions and guidance on if there is a certain  
order that images need to be retrieved in for future reference.
HSU:
1) Atlantic NAWIPS not displaying ASCAT wind barbs, only small circles.  
However, Pacific and Backup NAWIPS workstations are displaying ASCAT wind  
vectors just fine. TSB will investigate.

2) Still unable to launch GTWO on new Pacific dual-monitor NAWIPS  
workstation. Problem appears to be caused by gedit missing on that  
workstation so script hangs when it tries to open the text. TSB will work  
with NCO to get gedit back on that machine.

3) The email of SAB fixes did not work at 0000 or 0600 UTC. The fixes were  
also not on the Dvorak TC portion of the SAB web page. They can be found in  
the "bulletin" section of their webpage at:  
http://www.ssd.noaa.gov/PS/TROP/bulletins.html The fixes had to be manually  
entered into the ATCF by the HSU forecasters. HSU notified SAB of the  
problem and they will look into it. However, this issue is most likely  
related to the NCEP network problems yesterday. Please let TSB know if this  
continues to be a problem this week while NCO works to fix the networking.

4) Doug fixed the Convert program on the new RHEL7 ATCF workstations before  
he left on Tuesday.
Other items:
1) NCEP network outages yesterday have mostly been resolved, but there are  
still a few lingering issues. The NCEP/Centers network is working on a  
degraded state. We could still notice network issues today that could at  
times impact internet access, email, and non-SBN data flow issues going  
into either NAWIPS or AWIPS. Should HSU start advisories today, NHC will  
need to closely monitor the web to make sure everything is updating, since  
web services are also impacted by the network issues.

2) A reminder that today NCO will remove the /gpfs/hps/nhc links to  
/gpfs/hps3/nhc on the WCOSS Crays. TSB has made the necessary changes to  
our code, but please let TSB know if there is any difficulty connecting to  
or sending and retrieving files to/from WCOSS this week.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171004/a5762546/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list